what the equation of the line that contains the point (8,0)(0,2)
Found 2 solutions by puppylove12531, rfer:
Answer by puppylove12531(3)

(0-2)/(8-0)=-2/8=-1/4
-1/4 is the slope that is all I can help with

Answer by rfer(16322)

m=(2-0)/(0-8)
m=2/-8
m=-1/4
--
y=mx+b
y=-1/4x+2